# Packing subgraphs in regular graphs

### v1

@@ Abstract
An $H$-packing in a graph $G$ is a collection of pairwise vertex-disjoint copies of $H$ in $G$. We prove that for every $c>0$ and every bipartite graph $H$, any $\left\lfloor cn\right\rfloor$-regular graph $G$ admits an $H$-packing that covers all but a constant number of vertices. This resolves a problem posed by Kühn and Osthus in 2005. Moreover, our result is essentially tight: the conclusion fails if $G$ is not both regular and sufficiently dense; it is in general not possible to guarantee covering all vertices of $G$ by an $H$-packing, and if $H$ is not bipartite then $G$ need not contain any copies of $H$.
@@
Our proofs combine novel methods for balancing expanders and super-regular subgraphs with a number of powerful techniques including properties of robust expanders, the regularity lemma, and the blow-up lemma.
@@ 1 Introduction
Given two graphs $H$ and $G$, an $H$-packing (or an $H$-tiling) in $G$ is a collection of pairwise vertex-disjoint copies of $H$ in $G$. An $H$-packing in $G$ is called perfect if it covers all the vertices of $G$. Note that when $H$ consists of a single edge, an $H$-packing is a graph matching. Tutte’s theorem characterizes those graphs which have a perfect $H$-packing if $H$ is an edge, but no such characterisation is known for other connected graphs $H$. In fact, it is known [19] that the decision problem of whether a graph $G$ has a perfect $H$-packing is NP-complete if and only if $H$ has a component which contains at least three vertices. So it is natural to seek simple sufficient conditions that guarantee the existence of a perfect $H$-packing in a given graph.
A fundamental result in this area is the Hajnal–Szemerédi theorem [13] from 1970 which states that every graph whose order $n$ is divisible by $t$ and whose minimum degree is at least $(1-\frac{1}{t})n$ contains a perfect $K_{t}$-packing. For arbitrary graphs, Komlós, Sárközy and Szemerédi [22] showed that for every graph $H$, there exists a constant $C=C(H)$ such that every graph $G$ whose order $n$ is divisible by $|V(H)|$ and whose minimum degree is at least $(1-\frac{1}{\chi(H)})n+C$ contains a perfect $H$-packing. It turns out that there are graphs $H$ for which the bound on the minimum degree can be improved significantly i.e., one can often replace $\chi(H)$ by a smaller parameter called the critical chromatic number $\chi_{cr}(H)$ of $H$, which is defined as $\chi_{cr}(H)\coloneqq(\chi(H)-1)\frac{|H|}{|H|-\sigma(H)}$. Here $\sigma(H)$ denotes the minimum size of the smallest colour class in an optimal colouring of $H$. It is easy to see that $\chi(H)-1<\chi_{cr}(H)\leq\chi(H)$. In 2000, Komlós [20] showed that the critical chromatic number is the parameter that governs the existence of almost perfect packings in graphs of large minimum degree. Finally, in 2009, Kühn and Osthus [32] determined, up to an additive constant, the minimum degree of a graph $G$ that ensures the existence of a perfect $H$-packing in $G$, for every graph $H$. More precisely, they proved that for every graph $H$, either its critical chromatic number or its chromatic number is the parameter which governs the existence of perfect $H$-packings in graphs of large minimum degree (where the exact classification depends on a parameter called the highest common factor of $H$).
1.1 $H$-packings in dense regular graphs
In view of the above results, rather surprisingly, Kühn and Osthus [31] showed that if we restrict our attention to packings in regular graphs, then any linear bound on the degree guarantees an almost perfect $H$-packing.
@@ 1.1 $H$-packings in dense regular graphs
For every bipartite graph $H$ and every $0<c,\alpha\leq 1$, there exists $n_{0}=n_{0}(H,c,\alpha)$ such that every $d$-regular graph $G$ of order $n$, where $d\geq cn$ and $n\geq n_{0}$, has an $H$-packing that covers all but at most $\alpha n$ vertices of $G$.
It is easy to see that the restriction to (sufficiently) regular graphs $G$ in Theorem 1.1 is necessary. Indeed, if $G$ is, say, a complete bipartite graph $K_{n/4,3n/4}$ and $H$ is an edge, then it is clearly impossible to find an $H$-packing covering almost all the vertices of $G$. The restriction to bipartite graphs $H$ in Theorem 1.1 is also necessary because if $c\leq 1/2$, then $G$ could be a bipartite graph (in which case it cannot contain any subgraph $H$ which is non-bipartite).
This raises the natural question of whether the bound $\alpha n$ on the number of uncovered vertices in Theorem 1.1 can be lowered significantly to obtain an $H$-packing which is close to being perfect. Indeed, Kühn and Osthus [31] proposed the following general problem in 2005.
@@
Is it true that, for every bipartite graph $H$ and every $0<c\leq 1$, there is a constant $C=C(H,c)$ such that every $d$-regular graph $G$ of order $n$, where $d\geq cn$, has an $H$-packing that covers all but at most $C$ vertices of $G$?
Kühn and Osthus [31] resolved Problem 1.2 in the special case when the parts of $H$ have unequal sizes. More precisely, they showed the following.
@@
For every bipartite graph $H$ whose vertex classes have different size and for every $0<c\leq 1$, there is a constant $C=C(H,c)$ such that every $d$-regular graph $G$ of order $n$, where $d\geq cn$, has an $H$-packing which covers all but at most $C$ vertices of $G$.
@@
For every bipartite graph $H$ and every $0<c\leq 1$, there is a constant $C=C(H,c)$ such that every $d$-regular graph $G$ of order $n$, where $d\geq cn$, has an $H$-packing that covers all but at most $C$ vertices of $G$.
@@
As noted in [31], it is necessary to allow for uncovered vertices in Theorem˜1.4, and the upper bound $C$ on the number of uncovered vertices must depend on $c$ and $H$ (even if $n$ is divisible by $|V(H)|$), in sharp contrast to the result of Komlós, Sárközy and Szemerédi [22] mentioned earlier. Indeed, for instance, if $G$ is the vertex-disjoint union of cliques of order k|V(H)|1=d+1k|V(H)|-1=d+1 (for some positive integer $k$), then $G$ is $d$-regular, and it is easy to see that we must have at least $(|V(H)|-1)\cdot\frac{n}{d+1}$ uncovered vertices in any $H$-packing of $G$. Moreover, this example also shows that $d$ must be linear in $n$ in order for the number of uncovered vertices to be bounded by a constant $C$, so the requirement that the vertex degrees of $G$ are linear in $n$ in Theorem 1.4 is also necessary.
@@
It is easy to reduce Theorem˜1.4 to the special case when $H$ is a complete bipartite graph $K_{t,t}$. This follows, for example, from the fact that any bipartite graph $H$ can be perfectly packed into the complete bipartite graph $K_{|H|,|H|}$. To prove Theorem˜1.4 when $H=K_{t,t}$, we start by partitioning the vertex set of our dense regular graph $G$ into a small number of expanders using a structural decomposition result from the work of Gruslys and Letzter [11], which goes back to the work of Kühn, Lo, Osthus and Staden [27, 28] on Hamilton cycles in dense regular graphs. These expanders are, in turn, closely related to a powerful notion called robust expanders that was introduced by Kühn, Osthus and Treglown [34] and has since been used to prove several longstanding conjectures (see, e.g., [30, 29, 33]).
@@
One of the key contributions of our paper is a novel method for ‘balancing’ these expanders. This method enables us to obtain expanders that are either balanced bipartite or far from bipartite by removing a carefully constructed, small $K_{t,t}$-packing. Our main idea is to track a certain subgraph, which initially consists of the edges between expanders. We then sequentially move vertices between expanders and update the subgraph accordingly, ensuring that by the end, it has bounded maximum degree while still containing enough edges to construct a $K_{t,t}$-packing that balances the expanders (up to an additive constant, which is sufficient for our purposes). Note that while balancing expanders also plays a key role in [27, 28]—in resolving a problem of Bollobás and Häggkvist on Hamilton cycles in regular graphs—and in [11], which proves a conjecture of Magnant and Martin [39] for dense graphs, our approach differs significantly from the methods used in these works, so it is of independent interest. See Section 2.1 for further details on our balancing procedure.
@@ 1.2 Packing subdivisions in dense regular graphs
Quite a lot of research has established Conjecture 1.5 in several natural special cases. A result of Kelmans, Mubayi and Sudakov [18] proved the conjecture when $F$ is a tree, and in 2003, Alon [1] proved it when $F$ is a cycle. In 2005, Kühn and Osthus [31] observed that when $G$ is dense (i.e., $d=\Omega(n)$) the conjecture follows from their results on packings in dense regular graphs. Making significant progress towards the general case, Letzter, Methuku and Sudakov [36] applied their work on nearly-Hamilton cycles in sublinear expanders to show that the conjecture holds for all graphs $F$ when $G$ has degree at least polylogarithmic in $n$. Finally, very recently, Montgomery, Petrova, Ranganathan and Tan [42] resolved the conjecture in full.
For subdivisions of complete graphs $K_{t}$, Kühn and Osthus [31] went even further and showed that one can actually guarantee perfect packings in dense regular graphs for $t=4$ and $t=5$. More precisely, they showed that for every $0<c\leq 1$, there is a positive number $n_{0}=n_{0}(c)$ such that every $d$-regular graph $G$ of order $n$, with $d\geq cn$ and $n\geq n_{0}$, has a perfect $\operatorname{TK}_{t}$-packing for $t=4$ and $t=5$. By a result of Gruslys and Letzter [11], this also holds for $t=2$ and $t=3$.
@@
Kühn and Osthus’s result [31] actually guarantees a perfect $\operatorname{TK}_{t}$-packing, with $t\in\{4,5\}$, even when $G$ is only assumed to be almost regular. In contrast, for $t=3$ or $t\geq 6$, the restriction to regular graphs $G$ in Problem˜1.6 is necessary. Indeed, it is easy to see that the complete bipartite graph $K_{m,m+1}$ does not have a perfect $\operatorname{TK}_{t}$-packing if $t=3$ (note that in this case a $\operatorname{TK}_{3}$-packing corresponds to a cycle partition, i.e., a collection of pairwise vertex-disjoint cycles covering all vertices). Interestingly, Kühn and Osthus showed that $K_{m,m+1}$ also does not have a perfect $\operatorname{TK}_{t}$-packing if $t\geq 6$ (see Proposition 5.1 in [31]).
@@
To prove Theorem˜1.7, we begin by partitioning the vertex set of the dense regular graph $G$ into a small number of expanders, following the approach used in the proof of Theorem˜1.4. In this setting, however, we employ a lemma from [11] that provides a small linear forest $H$ that balances the expanders, ensuring that each expander contains either zero or two leaves of $H$. We then modify this linear forest so that at most one component of the new linear forest is associated with each expander, with both leaves of the component contained within that expander. Next, we construct a pair of small, vertex-disjoint $F$-subdivisions in each expander, while also ensuring that their union remains balanced. Finally, we absorb the components of $H$, along with any remaining uncovered vertices in each expander, into these $F$-subdivisions. This absorption step again relies on the Hamiltonicity of expanders and a robust connectivity property, which guarantees that any two vertices can be joined by a short path that avoids a small set of forbidden vertices. See Section 7 for a more detailed sketch of these ideas.
@@ 1.3 Organisation of the paper
In the next subsection, we present the notation used throughout the paper. In Section 2, we outline a detailed proof sketch of our main result, Theorem˜1.4, on almost perfect packings in dense regular graphs. In Section 3 we collect some tools and lemmas that are used throughout the paper. Key to our proof of Theorem˜1.4 are two lemmas – Lemma˜4.1 (for balancing expanders) and Lemma˜5.1 (for packing $K_{t,t}$’s in expanders). We prove Lemma˜4.1 in Section 4 and Lemma˜5.1 in Section 5, and complete the proof of Theorem˜1.4 in Section˜6. In Section 7, we prove our second result, Theorem˜1.7, which addresses perfect packings of subdivisions in dense regular graphs. A detailed sketch of its proof is given at the beginning of Section 7.
@@ 1.4 Notation
Using a slight abuse of notation, for a collection $\mathcal{K}$ of subgraphs in a graph $G$, we denote the set of vertices covered by these subgraphs as $V(\mathcal{K})$. Throughout this paper, we often omit floor and ceiling signs when dealing with large numbers, whenever they are not crucial.
2 Sketch of the proof of Theorem˜1.4
In this section, we present a sketch of the proof of our main result, Theorem˜1.4. At a high level, our approach begins by partitioning the vertex set of the regular graph $G$ into subsets with certain expansion properties, which we refer to as expanders. We then move some vertices between these expanders and remove a carefully constructed collection of $K_{t,t}$’s to ‘balance’ these expanders, using our first key lemma (Lemma˜4.1). Next, we find almost perfect $K_{t,t}$-packings in each of the resulting expanders with the help of our second key lemma (Lemma˜5.1). However, several significant challenges arise in implementing this approach, which we now address in detail.
Let $H$ be a bipartite graph, and let $0<c\leq 1$. Let $G$ be a $d$-regular graph of order $n$, where $d\geq cn$. Let $t=|V(H)|$. Recall that Theorem˜1.4 aims to find an $H$-packing that covers all but at most a constant number of vertices. It is easy to see that in order to find such a packing, it suffices to find a $K_{t,t}$-packing in $G$ covering all but at most $C$ vertices of $G$ for some constant $C=C(t,c)$ because $K_{t,t}$ has a perfect $H$-packing.
Our proof begins by partitioning the vertex set of the regular graph $G$ into a small number of subsets $Z_{1},\ldots,Z_{r}$ such that each induced subgraph $G[Z_{i}]$, for $i\in[r]$, has strong expansion properties, and the number of edges between different sets $Z_{i}$ is small. (See Lemma 3.2 for the precise statement.)

### v2

@@ Abstract
An $H$-packing in a graph $G$ is a collection of pairwise vertex-disjoint copies of $H$ in $G$. We prove that for every $c>0$ and every bipartite graph $H$, any $\left\lfloor cn\right\rfloor$-regular graph $G$ admits an $H$-packing that covers all but a constant number of vertices. This resolves a problem posed by Kühn and Osthus in 2005. Moreover, our result is essentially tight: the conclusion fails if $G$ is not both regular and sufficiently dense, it is in general not possible to guarantee covering all vertices of $G$ by an $H$-packing, and if $H$ is non-bipartite then $G$ need not contain any copies of $H$.
@@
Our proofs combine novel methods for balancing expanders and super-regular subgraphs with a number of powerful techniques including properties of robust expanders, regularity lemma, and blow-up lemma.
@@ 1 Introduction
Given two graphs $F$ and $G$, an $F$-packing (or an $F$-tiling) in $G$ is a collection of pairwise vertex-disjoint copies of $F$ in $G$. An $F$-packing in $G$ is called perfect if it covers all the vertices of $G$. Note that when $F$ consists of a single edge, an $F$-packing is a graph matching. Tutte’s theorem characterizes those graphs which have a perfect $F$-packing if $F$ is an edge, but no such characterisation is known for other connected graphs $F$. In fact, it is known [19] that the decision problem of whether a graph $G$ has a perfect $F$-packing is NP-complete if and only if $F$ has a component which contains at least three vertices. So it is natural to seek simple sufficient conditions that guarantee the existence of a perfect $F$-packing in a given graph.
A fundamental result in this area is the Hajnal–Szemerédi theorem [13] from 1970 which states that every graph whose order $n$ is divisible by $t$ and whose minimum degree is at least $(1-\frac{1}{t})n$ contains a perfect $K_{t}$-packing. For arbitrary graphs, Komlós, Sárközy and Szemerédi [22] showed that for every graph $F$, there exists a constant $C=C(F)$ such that every graph $G$ whose order $n$ is divisible by $|V(F)|$ and whose minimum degree is at least $(1-\frac{1}{\chi(F)})n+C$ contains a perfect $F$-packing. It turns out that there are graphs $F$ for which the bound on the minimum degree can be improved significantly i.e., one can often replace $\chi(F)$ by a smaller parameter called the critical chromatic number $\chi_{cr}(F)$ of $F$, which is defined as $\chi_{cr}(F)\coloneqq(\chi(F)-1)\frac{|F|}{|F|-\sigma(F)}$. Here $\sigma(F)$ denotes the minimum size of the smallest colour class in an optimal colouring of $F$. It is easy to see that $\chi(F)-1<\chi_{cr}(F)\leq\chi(F)$. In 2000, Komlós [20] showed that the critical chromatic number is the parameter that governs the existence of almost perfect packings in graphs of large minimum degree. Finally, in 2009, Kühn and Osthus [32] determined, up to an additive constant, the minimum degree of a graph $G$ that ensures the existence of a perfect $F$-packing in $G$, for every graph $F$. More precisely, they proved that for every graph $F$, either its critical chromatic number or its chromatic number is the parameter which governs the existence of perfect $F$-packings in graphs of large minimum degree (where the exact classification depends on a parameter called the highest common factor of $F$).
1.1 $F$-packings in dense regular graphs
In view of the above results, rather surprisingly, Kühn and Osthus [31] showed that if we restrict our attention to packings in regular graphs, then any linear bound on the degree guarantees an almost perfect $F$-packing.
@@ 1.1 $F$-packings in dense regular graphs
For every bipartite graph $F$ and every $0<c,\alpha\leq 1$, there exists $n_{0}=n_{0}(F,c,\alpha)$ such that every $d$-regular graph $G$ of order $n$, where $d\geq cn$ and $n\geq n_{0}$, has an $F$-packing that covers all but at most $\alpha n$ vertices of $G$.
It is easy to see that the restriction to (sufficiently) regular graphs $G$ in Theorem 1.1 is necessary. Indeed, if $G$ is, say, a complete bipartite graph $K_{n/4,3n/4}$ and $F$ is an edge, then it is clearly impossible to find an $F$-packing covering almost all the vertices of $G$. The restriction to bipartite graphs $F$ in Theorem 1.1 is also necessary because if $c\leq 1/2$, then $G$ could be a bipartite graph (in which case it cannot contain any subgraph $F$ which is non-bipartite).
This raises the natural question of whether the bound $\alpha n$ on the number of uncovered vertices in Theorem 1.1 can be lowered significantly to obtain an $F$-packing which is close to being perfect. Indeed, Kühn and Osthus [31] proposed the following general problem in 2005.
@@
Is it true that, for every bipartite graph $F$ and every $0<c\leq 1$, there is a constant $C=C(F,c)$ such that every $d$-regular graph $G$ of order $n$, where $d\geq cn$, has an $F$-packing that covers all but at most $C$ vertices of $G$?
Kühn and Osthus [31] resolved Problem 1.2 in the special case when the parts of $F$ have unequal sizes. More precisely, they showed the following.
@@
For every bipartite graph $F$ whose vertex classes have different size and for every $0<c\leq 1$, there is a constant $C=C(F,c)$ such that every $d$-regular graph $G$ of order $n$, where $d\geq cn$, has an $F$-packing which covers all but at most $C$ vertices of $G$.
@@
For every bipartite graph $F$ and every $0<c\leq 1$, there is a constant $C=C(F,c)$ such that every $d$-regular graph $G$ of order $n$, where $d\geq cn$, has an $F$-packing that covers all but at most $C$ vertices of $G$.
@@
As noted in [31], it is necessary to allow for uncovered vertices in Theorem 1.4, and the upper bound $C$ on the number of uncovered vertices must depend on $c$ and $F$ (even if $n$ is divisible by $|V(F)|$), in sharp contrast to the result of Komlós, Sárközy and Szemerédi [22] mentioned earlier. Indeed, for instance, if $G$ is the vertex-disjoint union of cliques of order k|V(F)|1=d+1k|V(F)|-1=d+1 (for some positive integer $k$), then $G$ is $d$-regular, and it is easy to see that we must have at least $(|V(F)|-1)\cdot\frac{n}{d+1}$ uncovered vertices in any $F$-packing of $G$. Moreover, this example also shows that $d$ must be linear in $n$ in order for the number of uncovered vertices to be bounded by a constant $C$, so the requirement that the vertex degrees of $G$ are linear in $n$ in Theorem 1.4 is also necessary.
@@
It is easy to reduce Theorem 1.4 to the special case when $F$ is a complete bipartite graph $K_{t,t}$. This follows, for example, from the fact that any bipartite graph $F$ can be perfectly packed into the complete bipartite graph $K_{|F|,|F|}$. To prove Theorem 1.4 when $F=K_{t,t}$, we start by partitioning the vertex set of our dense regular graph $G$ into a small number of expanders using a structural decomposition result from the work of Gruslys and Letzter [11], which goes back to the work of Kühn, Lo, Osthus and Staden [27, 28] on Hamilton cycles in dense regular graphs. These expanders are, in turn, closely related to a powerful notion called robust expanders that was introduced by Kühn, Osthus and Treglown [34] and has since been used to prove several longstanding conjectures (see, e.g., [30, 29, 33]).
@@
One of the key contributions of our paper is a novel method for ‘balancing’ these expanders. By this, we mean transforming expanders that are bipartite (or close to being bipartite) with imbalanced part sizes into balanced bipartite expanders. This method thus enables us to obtain expanders that are either balanced bipartite or far from bipartite by removing a carefully constructed, small $K_{t,t}$-packing. Our main idea is to track a certain subgraph, which initially consists of the edges between expanders. We then sequentially move vertices between expanders and update the subgraph accordingly, ensuring that by the end, it has bounded maximum degree while still containing enough edges to admit a $K_{t,t}$-packing that balances the expanders (up to an additive constant, which is sufficient for our purposes). Note that while balancing expanders also plays a key role in [27, 28]—in resolving a problem of Bollobás and Häggkvist on Hamilton cycles in regular graphs—and in [11], which proves a conjecture of Magnant and Martin [39] for dense graphs, our approach differs significantly from the methods used in these works, so it is of independent interest. See Section 2.1 for further details on our balancing procedure.
@@ 1.2 Packing subdivisions in dense regular graphs
Quite a lot of research has established Conjecture 1.5 in several natural special cases. The conjecture follows from a result of Kelmans, Mubayi and Sudakov [18] when $F$ is a tree, and in 2003, Alon [1] proved it when $F$ is a cycle. In 2005, Kühn and Osthus [31] observed that when $G$ is dense (i.e., $d=\Omega(n)$) the conjecture follows from their results on packings in dense regular graphs. Making significant progress towards the general case, Letzter, Methuku and Sudakov [36] applied their work on nearly-Hamilton cycles in sublinear expanders to show that the conjecture holds for all graphs $F$ when $G$ has degree at least polylogarithmic in $n$. Finally, very recently, Montgomery, Petrova, Ranganathan and Tan [42] resolved the conjecture in full.
For subdivisions of complete graphs $K_{t}$, Kühn and Osthus [31] went even further and showed that one can actually guarantee perfect packings in dense regular graphs for $t=4$ and $t=5$. More precisely, they showed that for every $0<c\leq 1$, there is a positive number $n_{0}=n_{0}(c)$ such that every $d$-regular graph $G$ of order $n$, with $d\geq cn$ and $n\geq n_{0}$, has a perfect $\operatorname{TK}_{t}$-packing for $t=4$ and $t=5$. By a result of Gruslys and Letzter [11], it follows that this also holds for $t=2$ and $t=3$.
@@
Kühn and Osthus’s result [31] actually guarantees a perfect $\operatorname{TK}_{t}$-packing, with $t\in\{4,5\}$, even when $G$ is only assumed to be almost regular. In contrast, for $t=3$ or $t\geq 6$, the restriction to regular graphs $G$ in Problem 1.6 is necessary. Indeed, it is easy to see that the complete bipartite graph $K_{m,m+1}$ does not have a perfect $\operatorname{TK}_{t}$-packing if $t=3$ (note that in this case a $\operatorname{TK}_{3}$-packing corresponds to a cycle partition, i.e., a collection of pairwise vertex-disjoint cycles covering all vertices). Interestingly, Kühn and Osthus showed that $K_{m,m+1}$ also does not have a perfect $\operatorname{TK}_{t}$-packing if $t\geq 6$ (see Proposition 5.1 in [31]).
@@
To prove Theorem 1.7, we begin by partitioning the vertex set of the dense regular graph $G$ into a small number of expanders, following the approach used in the proof of Theorem 1.4. In this setting, however, we employ a lemma from [11] that provides a small linear forest $H$ that balances the expanders, ensuring that each expander contains either zero or two leaves of $H$. We then modify this linear forest so that at most one component of the new linear forest is associated with each expander, with both leaves of the component contained within that expander. Next, we construct a pair of small, vertex-disjoint $F$-subdivisions in each expander, while also ensuring that their union remains balanced. Finally, we absorb the components of $H$, along with any remaining uncovered vertices in each expander, into these $F$-subdivisions. This absorption step again relies on the Hamiltonicity of expanders and a robust connectivity property, which guarantees that any two vertices can be joined by a short path that avoids a small set of forbidden vertices. See Section 7 for a more detailed sketch of these ideas.
@@ 1.3 Organisation of the paper
In the next subsection, we present the notation used throughout the paper. In Section 2, we outline a detailed proof sketch of our main result, Theorem 1.4, on almost perfect packings in dense regular graphs. In Section 3 we collect some tools and lemmas that are used throughout the paper. Key to our proof of Theorem 1.4 are two lemmas – Lemma 4.1 (for balancing expanders) and Lemma 5.1 (for packing $K_{t,t}$’s in expanders). We prove Lemma 4.1 in Section 4 and Lemma 5.1 in Section 5, and complete the proof of Theorem 1.4 in Section 6. In Section 7, we prove our second result, Theorem 1.7, which addresses perfect packings of subdivisions in dense regular graphs. A detailed sketch of its proof is given at the beginning of Section 7.
@@ 1.4 Notation
Using a slight abuse of notation, for a collection $\mathcal{K}$ of subgraphs in a graph $G$, we denote the set of vertices covered by these subgraphs as $V(\mathcal{K})$. Throughout this paper, we often omit floor and ceiling signs when dealing with large numbers whenever they are not crucial.
2 Sketch of the proof of Theorem 1.4
In this section, we present a sketch of the proof of our main result, Theorem 1.4. At a high level, our approach begins by partitioning the vertex set of the regular graph $G$ into subsets with certain expansion properties, which we refer to as expanders. We then move some vertices between these expanders and remove a carefully constructed collection of $K_{t,t}$’s to ‘balance’ these expanders, using our first key lemma (Lemma 4.1). More precisely, Lemma 4.1 allows us to transform expanders that are bipartite (or close to bipartite) with imbalanced part sizes into balanced bipartite expanders. Next, we find almost perfect $K_{t,t}$-packings in each of the resulting expanders with the help of our second key lemma (Lemma 5.1). Several significant challenges arise in implementing this approach, which we now address in detail.
Let $F$ be a bipartite graph, and let $0<c\leq 1$. Let $G$ be a $d$-regular graph of order $n$, where $d\geq cn$. Let $t=|V(F)|$. Recall that Theorem 1.4 aims to find an $F$-packing that covers all but at most a constant number of vertices of $G$. It is easy to see that in order to find such a packing, it suffices to find a $K_{t,t}$-packing in $G$ covering all but at most $C$ vertices of $G$ for some constant $C=C(t,c)$, because $K_{t,t}$ has a perfect $F$-packing.
Our proof begins by partitioning the vertex set of the regular graph $G$ into a small number of subsets $Z_{1},\ldots,Z_{r}$ (for some $r\leq\lceil 1/c\rceil$) such that, for each $i\in[r]$, the induced subgraph $G[Z_{i}]$ has strong expansion properties, is either almost bipartite or far from bipartite, and the number of edges of $G$ between different sets $Z_{i}$ is small. (See Lemma 3.2 for the precise statement.)